Can you help me with this problem. Thank you in advance...
find two numbers whose sum is 36 and such that one is 3 times as large as the other?
x=
y=

Answer by stanbon(75887)

find two numbers whose sum is 36 and such that one is 3 times as large as the other?
x+y = 36
x = 3y
--------------
Substitute to solve for "y":
3y + y = 36
4y = 36
y = 9
---------
Substitute to solve for "x":
x = 3y= 3*9 = 27
======================
Cheers,
Stan H.